Two sisters have ages that are consecutive even integers. The product of their ages is 120. Find the age of the sisters by writing and solving an algebraic equation. Make sure to define your variables.
Answer by Alan3354(69443)

Two sisters have ages that are consecutive even integers. The product of their ages is 120. Find the age of the sisters by writing and solving an algebraic equation. Make sure to define your variables.
---------
Find a pair of even integers with a product of 120 and a difference of 2.
--------
2*60 NG
4*30 NG
6*20 NG
10*12 That's it.
---------------------
n = younger sister's age.
n*(n+2) = 120
n^2 + 2n - 120 = 0
Find a pair of even integers with a product of 120 and a difference of 2. - Same as above.
2*60 NG
4*30 NG
6*20 NG
10*12 That's it.
---
(n - 10)*(n + 12) = 0
n = 10
Ages are 10 & 12.
A lot of extra steps for nothing.
---
Ignore the negative solution.